OpenStudy (anonymous):

Please help!!!!?:(The function b(x) = 2x − 5 determines how many bags of dog food need to be purchased for an animal shelter, where x is the number of dogs at the shelter. The shelter manager uses m(b(x)) to find the amount of money to bring for the dog food purchase. The function m(x) = 3x + 1. Solve for how much money to bring when there are 10 dogs in the shelter.

ganeshie8 (ganeshie8):

The shelter manager uses m(b(x))

ganeshie8 (ganeshie8):

use that function. put x = 10

OpenStudy (anonymous):

= 15

ganeshie8 (ganeshie8):

b(x) = 2x − 5 m(x) = 3x + 1 m(b(10)) = ?

ganeshie8 (ganeshie8):

yes, b(10) = 15

ganeshie8 (ganeshie8):

m(b(10) = m(15) = ?

ganeshie8 (ganeshie8):

find m(15)

ganeshie8 (ganeshie8):

m(x) = 3x + 1 put x = 15

OpenStudy (anonymous):

150

ganeshie8 (ganeshie8):

careful

OpenStudy (anonymous):

151

ganeshie8 (ganeshie8):

m(x) = 3x + 1 put x = 15 m(15) = 3(15) + 1 = ?

OpenStudy (anonymous):

46 my bad

ganeshie8 (ganeshie8):

m(b(10)) = m(15) = 46 so it takes 46$ to feed 10 dogs